# Memory-Augmented LLM Personalization with Short- and Long-Term Memory Coordination

### v1

@@ Abstract
Large Language Models (LLMs), such as GPT3.5, have exhibited remarkable proficiency in comprehending and generating natural language. However, their unpersonalized generation paradigm may result in suboptimal user-specific outcomes. Typically, users converse differently based on their knowledge and preferences. This necessitates the task of enhancing user-oriented LLM which remains unexplored. While one can fully train an LLM for this objective, the resource consumption is unaffordable. Prior research has explored memory-based methods to store and retrieve knowledge to enhance generation without retraining for new queries. However, we contend that a mere memory module is inadequate to comprehend a user’s preference, and fully training an LLM can be excessively costly. In this study, we propose a novel computational bionic memory mechanism, equipped with a parameter-efficient fine-tuning schema, to personalize LLMs. Our extensive experimental results demonstrate the effectiveness and superiority of the proposed approach. To encourage further research into this area, we are releasing a new conversation dataset generated entirely by LLM based on an open-source medical corpus, as well as our implementation code [^1] Github link will be placed here.
Introduction
The potential of large language models (LLMs) to understand and generate natural language is undeniable (Brown et al. 2020; Chowdhery et al. 2022; Touvron et al. 2023), while there is an untapped opportunity to explore how LLMs could be customised to provide personalized interactions with users, allowing them to receive tailored responses that best suit their individual needs (Bender et al. 2021). Efforts have been made to design proper prompts for steering LLMs to enhance output. For example, by memorizing previous mistakes and user feedback, given a new query, a similarity-based retriever can be leveraged to preemptively recognize and rectify LLM errors(Dalvi, Tafjord, and Clark 2022; Madaan et al. 2022; Lewis et al. 2020). However, this paradigm poses us two challenges: Firstly, most existing memory designs are dictionary-based (Madaan et al. 2022; Lewis et al. 2020) (i.e. key-value form where key is the previous mistake, value is the corresponding user-feedback) which can be inflexible and rely heavily on the power of retriever. For example, as the conversation goes, the size of memory in this setting would increase uncontrollably which will further impact the effectiveness and accuracy of the retriever; Secondly, such paradigm, without retraining, can barely provide users with personalized and engaging experience. For instance, in the healthcare context, a diabetes patient who prefers concise and straightforward medical advice won’t expect detailed glucose test explanations from a doctor while others who prefer fully elaborated responses may want to know as much as possible about the disease (e.g., causes etc.). To this end, being aware of user preference and user-relevant information can be crucial for enhancing user experience and remains understudied. In this paper, we propose a novel framework that enhances personalized LLM learning through the augmentation of memory and parameters-efficient fine-tuning schema.
@@ Introduction
Dictionary-based memory is not pliable due to its intricate structure and thus efforts can only be made in strengthening retrievers. Despite the improvements made by retrievers like semantic-similarity based and distance-closest based(Madaan et al. 2022), we argue that the memory structure should be ameliorated to accommodate diverse information. For example, as Figure 1 depicted, doctors would recognize patients’ essential information after years of engaging in diagnostic conversations. Among such information, however, some would be memorized in a long-last way such as chronic diseases, while some are not and those memories can be used for enhancing the responses in the future dialogues. Unfortunately, rare efforts have been made to address this challenge. Neuroscientists have revealed that real-world memory mechanism works in two processes, one can be characterized as automatic and unconscious while the other one is effortful and analytical(Kahneman 2011). For example, acquiring a new skill initially can be active and analytical, but as one’s proficiency increases, it becomes more intuitive. This is referred as Dual-process theory and correspondingly, memory can be defined as three types: working memory, short-term memory (STM) and long-term memory (LTM). Working memory is responsible for filtering and buffering information, STM holds knowledge for a short period, while LTM stores knowledge for a longer duration (Roediger and McDermott 1995). Drawing inspirations from this, we propose a novel Dual-Process enhanced Memory (DPeM) mechanism of which three types of memory cooperate smoothly under the guidance of dual-process schema and thus can provide LLM more useful knowledge from both user-specific and common-sense aspects.
@@
Furthermore, existing works in personalized LLMs primarily focus on designing comprehensive prompts that enable LLMs to generate contextually relevant responses aligned with user-specific dialogue preferences (e.g., preferring concise responses) (Wang et al. 2023; Wu et al. 2023; Wang, Ustun, and Calmon 2019). However, these methods often yield inferior performance compared to fine-tuning approaches and are susceptible to the exact formatting of the prompts (e.g., wording and ordering) (Liu et al. 2022a). Another approach proposed by Salemi et al. (2023) involves incorporating user profiles during the pretraining stage, enabling LLMs to possess user-specific knowledge for downstream tasks. Nevertheless, training fully personalized LLMs for individuals can be economically unviable. To this end, we opt to the utilization of parameter-efficient fine-tuning (PEFT) techniques to develop user-oriented LLMs with reduced time and resource consumption. Specifically, PEFT focuses on updating a small subset of parameters, ensuring that the trained LLM achieves promising performance on new tasks while minimizing computational costs. PEFT can enable LLM to learn to respond to users while being aware of their specific needs in a computationally and resource-efficient way.
@@
In tandem, we propose a novel memory mechanism inspired by neuroscience, and along with a PEFT training strategy to achieve LLM personalization. Our approach moves away from the simplistic dictionary-based structure, instead closely resembling real-world memory processes. The key contributions of our work are as follows:
@@
$\bullet$ We propose a novel DPeM mechanism that closely resembles real-world memory processes which lead to a more efficient augmentation for LLM.
@@
$\bullet$ We propose MaLP, a unified frame based on DPeM and PEFT which aims to promote the LLM’s quality of responding by catering to user-specific needs.
$\bullet$ Our proposed method outperforms several baselines on three tasks and the conducted human evaluation further validate the effectiveness and superiority of MaLP.
@@
$\bullet$ We introduce a new dialogue dataset in the medical domain that incorporates user preferences and historical records. This dataset offers a unique opportunity to explore personalized LLM models from a fresh perspective.
@@
Methodology
Preliminary Definition
DPeM Mechanism
The key novelty that differentiates our work with previous efforts is that we turns into excavating the improvements regarding the intricate memory structure instead of studying solely on retriever. Drawing inspirations from Kahneman (2011), we aim to design a memory mechanism that closely resembles real-world memory processes. To achieve this, we propose a dual-process (Rehearsal Process, Executive Process) enhanced procedure which consists of three steps - Learning, Summarizing and Memorizing. The Rehearsal Process involves learning information from $\mathcal{D}$, which is then stored in working memory. The working memory is refreshed iteratively based on the dialogue’s content of current iteration—this is the summarizing step. The two-step rehearsal process is facilitated by a coordinator with powerful natural language understanding abilities. Next, the information stored in working memory is evaluated to determine whether it needs to be stored in Short-Term Memory (STM) or Long-Term Memory (LTM) based on the frequency of access by the Executive Process. This dual-process is illustrated in Figure 2 by the green-box and the three-colored lines. The detailed workings of DPeM are as follows:
Memory $\mathcal{M}$:
As $\mathcal{M}$ does not consist of a single type of memory. Instead, it comprises different types of memory that store and access information in their own way, while working together for better knowledge management through dual-process. Intuitively, there will be a vast amount of information that needs to be registered when learning something new and not all those information will be stored directly and entirely into memory. Instead, a working memory acts as a buffer memory to register and filter information so that only the relevant information enters STM, while the rest is dropped. Short-term memory refers to a limited space that holds a small amount of knowledge in an active, quickly accessible state. Long-term memory stores knowledge transited from STM for a longer period. Being aware of the differences in information storage and access, our memory mechanism enhances LLM by incorporating knowledge from both user-specific and common-sense perspectives. All the three types of memory are in the form of a growing table which support different operations but work collaboratively as shown in Table 1.
Rehearsal Process
refers to the process of obtaining and refreshing information so that the relevant information can be filtered and stored as knowledge into memory. Learning aims at gathering information from the dialogue’s content of the current iteration. Inspired by how humans take notes when learning something new, we leverage $\mathcal{C}$, which takes all the necessary notes for each iteration’s dialogue. Formally, by passing $d_{i}\in\mathcal{D}$ to $\mathcal{C}$ iteratively, we obtain some notes denoted as $nts=\mathcal{C}(d_{i})$ which will be recorded into working memory. Typically, not all the notes will be practical which makes the summarizing step indispensable within the DPeM mechanism. Summarizing steps further by purifying the relevant notes from $nts$ and pass them into STM. Specifically, the coordinator will determine if the $nt\in nts$ is relevant or not and store the useful $nt^{+}$ as knowledge item by item and the stored knowledge is delivered to STM and is denoted as $K=[k_{0},...,k_{m}]$ where $k_{i}=nt_{i}^{+}$.
@@
Executive Process
mainly focuses on manipulating and memorizing the knowledge produced from rehearsal process. The main objective of DPeM is to process and store information based on its importance level and user-specific needs, an aspect that previous works have paid little attention to. Specifically, the filtered knowledge $k$ will firstly be categorized as two types: Common-sense Knowledge, User-Specific Knowledge and then be converted into the STM in the form of key (type) - value ($k_{i}$) pair. As the learning iteration progresses, a flag table $ft$ is used to keep track of the frequency of appearance for each $k_{i}$. When the frequency reaches a predetermined threshold $\theta$, the $k_{i}$ is transferred to LTM. Notably, STM is refreshed periodically after certain rounds (working memory is refreshed after each iteration) while LTM typically only accepts new $k_{i}$ entries. The final memory structure consists of three parts: Working Memory, STM, and LTM. Working Memory serves as a buffer for storing newly detected information, STM stores relevant and recent knowledge, and LTM provides longer-term access to frequently visited knowledge from STM. Through the collaboration of these three types of memory, along with the dual-process approach, DPeM provides a more powerful memory system to further support personalized LLM.
MaLP Frame
Memory Generation
@@
where $nt_{i}=\mathcal{C}(d_{i})$, $k_{j}=nt_{i}^{+}$, $k_{f}$ denotes frequently visited $k_{j}$ from $\mathcal{M}_{STM}$. The comparison among these three types of memory can be seen in Table 1.
Memory Utilization
However, relying solely on memory for achieving personalized LLMs still poses challenges, as the quality of generated responses ultimately depends on the understanding and generation ability of the LLM, even with memory-augmented prompts. Therefore, fine-tuning the LLM to cater to user-specific needs naturally becomes an option for enhancing LLM personalization. However, traditional fine-tuning approaches often demand significant computational and data resources, whereas our aim is to optimize the LLM’s response generation in a user-friendly manner by leveraging previous dialogues. In this regard, PEFT methods (Li and Liang 2021; Liu et al. 2022b, c) offer a solution by achieving this objective with low resource consumption.
To tune the LLaMA model with a user’s previous dialogues and enable it to generate user-favorable responses, we employ the Low-Rank Adaption (LoRA) technique (Hu et al. 2021). With LoRA, we update a given pre-trained weight matrix $W_{\Phi}\in\mathbb{R}^{d\times k}$ of LLaMA by incorporating a low-rank decomposition $W_{\Phi}+\Delta W=W_{\Phi}+BA$. Here, $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During the fine-tuning process, we randomly select a set number of layers to implement LoRA, where $A$ and $B$ are trainable while $W_{\Phi}$ remains frozen. This approach allows us to target user preferences and adapt the LLaMA model accordingly.
Once all iterations are completed, we acquire a LoRA-tuned LLaMA model along with a latent memory that caters to user-specific needs. Typically, when a new query $x$ is received, the responding process is presented as:
@@
where $p$ is the prompt retrieved from $\mathcal{M}$, $Retriever$ is a function that can retrieve knowledge from $\mathcal{M}$ in terms of query $x$ and $\hat{\Phi}$ is LoRA-tuned LLaMA. The utilization process is denoted in orange lines as can be seen in Figure 2.
Components
@@
In the MaLP framework, several key components actively engage in memory generation and utilization to ensure efficient and collaborative operations:
@@
Coordinator $\mathcal{C}$: $\mathcal{C}$ plays a pivotal role in the learning and summarizing stage which involves deriving information from dialogue contexts and purifying knowledge from learned information (Xu et al. 2023). Thus we resort to a powerful tool (e.g., ChatGPT) that is capable of understanding the long dialogue and performing summarization and judgement(Xu et al. 2023).

### v3

# LLM-based Medical Assistant Personalization with Short- and Long-Term Memory Coordination

@@ Abstract
Large Language Models (LLMs), such as GPT3.5, have exhibited remarkable proficiency in comprehending and generating natural language. On the other hand, medical assistants hold the potential to offer substantial benefits for individuals. However, the exploration of LLM-based personalized medical assistant remains relatively scarce. Typically, patients converse differently based on their background and preferences which necessitates the task of enhancing user-oriented medical assistant. While one can fully train an LLM for this objective, the resource consumption is unaffordable. Prior research has explored memory-based methods to enhance the response with aware of previous mistakes for new queries during a dialogue session. We contend that a mere memory module is inadequate and fully training an LLM can be excessively costly. In this study, we propose a novel computational bionic memory mechanism, equipped with a parameter-efficient fine-tuning (PEFT) schema, to personalize medical assistants. To encourage further research into this area, we are releasing a new conversation dataset generated based on an open-source medical corpus and our implementation code [^1] https://github.com/MatthewKKai/MaLP.
1 Introduction
@@
The potential of large language models to understand and generate natural language is undeniable (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023), while there is an untapped opportunity to explore how LLMs could be customised to provide personalized medical advice with patients, allowing them to receive tailored responses that best suit their individual needs (Bender and Koller, 2020). For example, as depicted in Figure 1, medical practitioners can discern vital patient information through ongoing diagnostic conversations. Consequently, responses to identical queries may differ based on individual patient nuances, highlighting the imperative need for personalized medical assistants leveraging LLM. Efforts have been made to obtain proper prompts for steering LLMs to enhance outputs. For example, by memorizing previous mistakes and user feedback, given a new query, a similarity-based retriever can be leveraged to preemptively recognize and rectify LLM errors(Dalvi et al., 2022; Madaan et al., 2022; Lewis et al., 2020). However, this paradigm poses us two challenges: Firstly, most existing memory designs are dictionary-based Madaan et al. (2022); Lewis et al. (2020) (i.e. key-value form where key is the previous mistake, value is the corresponding user-feedback) which can be inflexible and rely heavily on the power of retriever. Secondly, such paradigm, without retraining, can barely provide users with personalized and engaging experience. For instance, a diabetes patient who prefers concise and straightforward medical advice won’t expect detailed glucose test explanations from a doctor while others who prefer fully elaborated responses may want to know as much as possible about the disease (e.g., causes etc.). To this end, how to process patient-relevant information properly and being aware of their preference can be crucial for enhancing patients’ experience and remains understudied. In this paper, we propose a novel memory mechanism along with a PEFT schema to enhance LLM-based medical assistant personalization.
@@
Dictionary-based memory is not pliable due to its intricate structure and thus efforts can only be made in strengthening retrievers. Despite the improvements made by retrievers like semantic-similarity based and distance-closest basedMadaan et al. (2022), we argue that the memory structure should be ameliorated to accommodate diverse information. Unfortunately, rare efforts have been made to address this challenge. Neuroscientists have revealed that real-world memory mechanism works in two processes, one can be characterized as automatic and unconscious while the other one is effortful and analytical(Kahneman, 2011). For example, acquiring a new skill initially can be active and analytical, but as one’s proficiency increases, it becomes more intuitive. This is referred as Dual-process theory and correspondingly, memory can be defined as three types: working memory, short-term memory (STM) and long-term memory (LTM). Working memory is responsible for filtering and buffering information, STM holds knowledge for a short period, while LTM stores knowledge for a longer duration (Roediger and McDermott, 1995). Drawing inspirations from this, we propose a novel Dual-Process enhanced Memory (DPeM) mechanism of which three types of memory cooperate smoothly under the guidance of dual-process schema and thus can provide LLM more useful knowledge from both user-specific and common-sense aspects.
@@
Furthermore, existing works in personalized LLMs primarily focus on designing comprehensive prompts that enable LLMs to generate contextually relevant responses aligned with user-specific dialogue preferences (e.g., preferring concise responses) (Wang et al., 2023; Wu et al., 2023; Wang et al., 2019). However, these methods often yield inferior performance compared to fine-tuning approaches and are susceptible to the exact formatting of the prompts (e.g., wording and ordering) (Liu et al., 2022a). Another approach proposed by Salemi et al. (2023) involves incorporating user profiles during the pretraining stage, enabling LLMs to possess user-specific knowledge for downstream tasks. Nevertheless, training fully personalized LLMs for individuals can be economically unviable. To this end, we embrace the utilization of PEFT which focuses on updating a small subset of parameters, ensuring that the trained LLM achieves promising performance on new tasks while minimizing computational costs, to develop user-oriented LLMs with reduced time and resource consumption.
@@
In tandem, we propose a novel memory mechanism inspired by neuroscience, and along with a PEFT training strategy to achieve LLM-based medical assistant personalization. The key contributions of our work are as follows:
@@
$\bullet$ We propose a novel DPeM mechanism that closely resembles real-world memory processes which lead to a relatively 7% improvement against existing memory structure.
@@
$\bullet$ We propose MaLP, a unified frame based on DPeM and PEFT which promotes the response’s quality by catering to user-specific needs.
@@
$\bullet$ We introduce a new medical dialogue dataset that incorporates user preferences and historical records. This dataset offers a unique perspective to explore personalized medical assistants.
2 Methodology
2.1 Preliminary Definition
2.2 Medical Knowledge Adaptation
To help the LLM provide better responses, we propose to first incorporate more medical knowledge via a domain adapter(Zhang et al., 2023). The adapter architecture consists of a down-projection layer, a non-linearity function (e.g., ReLU(Agarap, 2018)), and an up-projection (e.g., a fully connected network). Note that all parameters, except those pertaining to the domain adapter, remain frozen. However, directly using such domain adapter will lead to the catastrophic forgetting problem(Gururangan et al., 2020). This phenomenon entails the risk that the LLM may lose its inherent capabilities after training on domain-specific knowledge using an adapter. To solve this, we propose leveraging a sample loss, which gauges the output disparity before and after the knowledge adaptation process for the same query. Given a medical text with K masked tokens, the knowledge loss can be $\mathcal{L}_{K}=-\frac{1}{K}\sum_{i=1}^{M}log\ p(m_{i})$ where $p(m_{i})$ is the probability of generating $m_{i}$, and the sample loss can be defined as $\mathcal{L}_{S}=||V_{o},V_{k}||_{2}^{2}$, where the $V_{o}$ is the vector representation of tokens from the original layer and $V_{k}$ is the vector representation of tokens from the layer that installed the adapter. The overall knowledge adaptation can be trained by simply adding those two losses. We use the trained LLM with external medical knowledge as the base LLM in the following steps.
2.3 DPeM Mechanism
The key novelty that differentiates our work with previous efforts is that we turns into excavating the improvements regarding the intricate memory structure instead of studying solely on retriever. Drawing inspirations from Kahneman (2011), we aim to design a memory mechanism that closely resembles real-world memory processes. To achieve this, we propose a dual-process (Rehearsal Process, Executive Process) enhanced procedure which consists of three steps - Learning, Summarizing and Memorizing. The Rehearsal Process involves learning information from $\mathcal{D}$, which is then stored in working memory. The working memory is refreshed iteratively based on the dialogue’s content of current iteration—this is the summarizing step. The two-step rehearsal process is facilitated by a coordinator with powerful natural language understanding abilities. Next, the information stored in working memory is evaluated to determine whether it needs to be stored in Short-Term Memory (STM) or Long-Term Memory (LTM) based on the frequency of access by the Executive Process. This dual-process is illustrated in Figure 2 by the green-box and the three-colored lines. The detailed memory structure and working flows of DPeM are depicted in the following sections.
@@ 2.3 DPeM Mechanism
2.3.1 Memory $\mathcal{M}$:
As prescribed, $\mathcal{M}$ does not consist of a single type of memory. Instead, it comprises different types of memory that store and access information in their own way, while working together for better knowledge management through dual-process. Intuitively, there will be a vast amount of information that needs to be registered when learning something new and not all those information will be stored directly and entirely into memory. Instead, a working memory acts as a buffer memory to register and filter information so that only the relevant information enters STM, while the rest is dropped. Short-term memory refers to a limited space that holds a small amount of knowledge in an active, quickly accessible state. Long-term memory stores knowledge transited from STM for a longer period. Being aware of the differences in information storage and access, our memory mechanism enhances LLM by incorporating knowledge from both user-specific and common-sense perspectives. All the three types of memory are in the form of a growing table which support different operations but work collaboratively as shown in Table 1.
2.3.2 Rehearsal Process
Rehearsal refers to the process of obtaining and refreshing information so that the relevant information can be filtered and stored as knowledge into memory. Learning aims at gathering information from the dialogue’s content of the current iteration. Inspired by how humans take notes when learning something new, we leverage $\mathcal{C}$, which takes all the necessary notes for each iteration’s dialogue. Formally, by passing $d_{i}\in\mathcal{D}$ to $\mathcal{C}$ iteratively, we obtain some notes denoted as $nts=\mathcal{C}(d_{i})$ which will be recorded into working memory. Typically, not all the notes will be practical which makes the summarizing step indispensable within the DPeM mechanism. Summarizing steps further by purifying the relevant notes from $nts$ and pass them into STM. Specifically, the coordinator will determine if the $nt\in nts$ is relevant or not and store the useful $nt^{+}$ as knowledge item by item and the stored knowledge is delivered to STM and is denoted as $K=[k_{0},...,k_{m}]$ where $k_{i}=nt_{i}^{+}$.
@@ 2.3.2 Rehearsal Process
2.3.3 Executive Process
Executive Process mainly focuses on and memorizing the knowledge produced from rehearsal process. The main objective of DPeM is to process and store information based on its importance level and user-specific needs, an aspect that previous works have paid little attention to. Specifically, the filtered knowledge $k$ will firstly be categorized as two types: Common-sense Knowledge, User-Specific Knowledge and then be converted into the STM in the form of key (type) - value ($k_{i}$) pair. As the learning iteration progresses, a flag table $ft$ is used to keep track of the frequency of appearance for each $k_{i}$. When the frequency reaches a predetermined threshold $\theta$, the $k_{i}$ is transferred to LTM. Notably, STM is refreshed periodically after certain rounds (working memory is refreshed after each iteration) while LTM typically only accepts new $k_{i}$ entries. The final memory structure consists of three parts: Working Memory, STM, and LTM. Working Memory serves as a buffer for storing newly detected information, STM stores relevant and recent knowledge, and LTM provides longer-term access to frequently visited knowledge from STM. Through the collaboration of these three types of memory, along with the dual-process approach, DPeM provides a more powerful memory system to further support personalized LLM.
2.4 MaLP Frame
2.4.1 Memory Generation
@@
where $nt_{i}=\mathcal{C}(d_{i})$, $k_{j}=nt_{i}^{+}$, $k_{f}$ denotes frequently visited $k_{j}$ from $\mathcal{M}_{STM}$. The comparison among these three types of memory can be seen in the Table 1.
2.4.2 Memory Utilization
However, relying solely on memory for achieving personalized LLMs still poses challenges, as the quality of generated responses ultimately depends on the understanding and generation ability of the LLM, even with memory-augmented prompts and pre-injected knowledge. Therefore, fine-tuning the LLM to cater to user-specific needs naturally becomes an option for enhancing LLM personalization. However, traditional fine-tuning approaches often demand significant computational and data resources, whereas our aim is to optimize the LLM’s response generation in a user-friendly manner by leveraging previous dialogues. In this regard, PEFT methods (Li and Liang, 2021; Liu et al., 2022b, c) offer a solution by achieving this objective with low resource consumption.
To tune the base LLM (e.g., LLaMA) with user’s previous dialogues and enable it to generate user-favorable responses, we employ the Low-Rank Adaption (LoRA) technique (Hu et al., 2021). With LoRA, we update a given pre-trained weight matrix $W_{\Phi}\in\mathbb{R}^{d\times k}$ of LLM by incorporating a low-rank decomposition $W_{\Phi}+\Delta W=W_{\Phi}+BA$. Here, $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During the fine-tuning process, we randomly select a set number of layers to implement LoRA, where $A$ and $B$ are trainable while $W_{\Phi}$ remains frozen. This allows us to target user preferences and adapt the LLM accordingly.
Once all iterations are completed, we acquire a LoRA-tuned LLM along with a latent memory that caters to user-specific needs. Typically, when a new query $x$ is received, the responding process is presented as:
@@ 2.4.2 Memory Utilization
where $p$ is the prompt retrieved from $\mathcal{M}$, $Retriever$ is a function that can retrieve knowledge from $\mathcal{M}$ in terms of query $x$ and $\hat{\Phi}$ is the LoRA-tuned LLM. The utilization process is denoted in orange lines as can be seen in Figure 2.
2.4.3 Components
@@
In the MaLP framework, several key components actively engage in memory generation and utilization to ensure efficient collaborations:
@@
Coordinator $\mathcal{C}$: $\mathcal{C}$ plays a pivotal role in the learning and summarizing stage which involves deriving information from dialogue contexts and purifying knowledge from learned information (Xu et al., 2023). Thus we resort to a powerful tool (e.g., ChatGPT) that is capable of understanding the long dialogue and performing summarization and judgement(Xu et al., 2023).